OpenStudy (anonymous):

If the midpoint between (8, y) and (-11, 6) is (-1.5, 5), find the value of y.

OpenStudy (anonymous):

(y1+y2)/2=5 (y1+6)/2=5 y1+6=10 y1=10-6 y1=4